[figure: 9.7 cm]
μSR STUDY ON CeRh_{2}Si_{2}

D. Andreica^{a}, A. Amato^{b}, F.N. Gygax^{c} and A. Schenck^{c} a Faculty of Physics, Babeş-Bolyai University, RO-3400 Cluj-Napoca, Romania b Laboratory for μSR, Paul Scherrer Institutr, 5232-Villgen-PSI, Switzerland c Institute for Particle Physics, ETH-Zürich, Switzerland

CeRh_{2}Si_{2} has two magnetic transition temperatures, T_{N1} and T_{N2}. Using the μSR technique we found the suppression - by the applied pressure - of the magnetic transition temperatures, in agreement with results obtained in macroscopic measurements [GR97]. However, the magnetic structure proposed from neutron scattering under pressure measurements [KA00], are at variance with our μSR results both below T_{N1} and T_{N2}. At temperatures between T_{N1} and T_{N2} the origin of one of the two contributions to the μSR signal cannot be explained by the proposed magnetic structure. Below T_{N2} our results point to a double domain magnetic structure, in contrast with the single domain 4-k magnetic structure proposed in [KA00]. Moreover, according to the neutron scattering results the ordered moment decreases with increasing pressure, while the μSR results reveal a pressure independent ordered moment, at least up to 5 kbar. These results indicate a well localized 4f moment for the Ce ions, in the low pressure range, in agreement with the large magnetic transition temperature of CeRh_{2}Si_{2}, T_{N1} = 36 K. The magnetic structure of CeRh_{2}Si_{2} remains a subject of debate.

[GR97] F. M. Grosche et al., Physica B 237-238 (1997) 197.

[KA00] S. Kawarazaki, et al, Phys. Rev. B 61 (2000) 4167.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : D. Andreica

Address for correspondence : Faculty of Physics Babeş-Bolyai University RO-3400 Cluj-Napoca ROMANIA

Email address : dandr@phys.ubbcluj.ro